I have no idea how to go about this problem...or where to start. Iknow this has to do with the distance, time, and rate process but I'm still not sure how to go about this problem or to solve it. Help me please :(
An airplane flies 1,050 km with the wind . In the same amount of time, it can fly 750 km against the wind. The speed of the plane in still air is 200 km/h. Find the speed of the wind.

Answer by sachi(548)

The speed of the plane in still air is 200 km/h.
let the speed of wind=x km/h
airplane flies 1,050 km with the wind
here the speed =200+x km/h
so the time= 1050/(200+x) hr
can fly 750 km against the wind
here the speed =200-x km/h
so the time= 750/(200-x) hr
the same amount of time is taken in both flights
so1050/(200+x)=750/(200-x)
or 7(200-x)=5(200+x)
or 1400-7x=1000+5x
or 12x=400
or x=the speed of the wind-400/12=33.33 km/h